# Revisiting Jahn–Teller Transitions in Correlated Oxides with Monte Carlo Modeling

### v2

@@ Abstract
Jahn–Teller (JT) distortions are a key driver of physical properties in many correlated oxide materials. Cooperative JT distortions, in which long-range orbital order reduces the symmetry of the average structure macroscopically, are common in JT-distorted materials at low temperatures. This long-range order will often melt on heating, via a transition to a high-temperature state without long-range orbital order. The nature of this transition has been observed to vary with different materials depending on crystal structure; in LaMnO3 the transition has generally been interpreted as order-disorder, whereas in layered nickelates AANiO2 (AA=Li,Na) there is a displacive transition. Alternatively, recent theoretical work has suggested that previous attributions of order-disorder may in fact be a consequence of phonon anharmonicity, rather than persistence of JT distortions, which would suggest that the displacive transition may be more common than currently believed. In this work, we run Monte Carlo simulations with a simple Hamiltonian which is modified to include terms dependent on the JT amplitude $\rho$, which is allowed to vary within the simulation via the Metropolis algorithm. Our simulations yield distributions of JT amplitudes consistent with displacive rather than order-disorder behaviour for both perovskites and layered nickelates, suggesting that displacive-like JT transitions may be more common than previously assumed in both perovskites and layered nickelates. We also find significant differences between the transition observed for perovskites compared with layered nickelates, which we attribute to differing extensivity of configurational entropy on the two lattices, showing the crucial role of lattice geometry in determining behaviour.
@@ I Introduction
Transition metal oxide octahedra with electronic configurational degeneracy are susceptible to a Jahn–Teller (JT) distortion Jahn and Teller (1937); Öpik and Pryce (1957); Longuet-Higgins et al. (1958); Kanamori (1960); Gehring and Gehring (1975); Halcrow (2013); Goodenough (1998), causing their shape to deviate from that of a regular octahedron. Such distortions most commonly manifest as a linear superposition of the $E_{g}(Q_{2},Q_{3})$ van Vleck modes Van Vleck (1939); Nagle-Cocco and Dutton (2024), where $Q_{2}$ and $Q_{3}$ are planar rhombic and tetragonal distortions, respectively, although the tetragonal $Q_{3}$ component tends to dominate. JT distortions have been observed in both molecules Halcrow (2013) and crystals Goodenough (1998). The JT distortion, and associated orbital ordering, is relevant to many phenomena including superconductivity in the cuprates Fil et al. (1992); Keller et al. (2008); Bussmann-Holder and Keller (2022), spin-orbit ordering Khomskii and Streltsov (2021), and ionic mobility Kim et al. (2015); Li et al. (2016), and can lead to structural transitions on cycling in battery electrode materials Kim et al. (2015); Li et al. (2016); Choi et al. (2019).
@@
Within solids, the axes of octahedral elongation can correlate via orbital ordering over macroscopic length-scales as a cooperative JT (cJT) distortion, typically manifesting as a reduction in unit cell symmetry Gehring and Gehring (1975); Goodenough (1998). For instance, the perovskites KMF3 (M=Cr,Cu) Zhou et al. (2011); Margadonna and Karotsis (2006), LaMnO3 Rodríguez-Carvajal et al. (1997), and [(CH3)2NH2]Cu(HCOO)3 Scatena et al. (2021) exhibit a cJT distortion with alternating MO6 (M=Cu,Cr,V,Mn) octahedra elongated along two orthogonal directions within the $xy$-plane, as shown in Figure 1(a). This ordering is generally stacked identically along the $c$-direction Rodríguez-Carvajal et al. (1997); Ren et al. (1998); Blake et al. (2009); Zhou et al. (2011). Interestingly, similar orbital ordering driven by $t_{2g}$ orbitals is seen in the vanadate perovskites AVO3 (A=Ho,La,Y) Blake et al. (2009); Ren et al. (1998); Blake et al. (2001, 2002), although on cooling to low temperatures, some (such as LaVO3 and YVO3) have been shown to exhibit out-of-phase stacking along the $c$-direction Bordet et al. (1993); Blake et al. (2001, 2002) [SI Figure S1]. The flexibility of the perovskite structure means that other kinds of ordering can also be exhibited; Ba0.5La0.5CoO3 exhibits a collinear elongational cJT ordering of the CoO6 octahedra when the La/Ba is disordered Fauth et al. (2002) but a compressive collinear ordering when the La/Ba is ordered Nakajima et al. (2005) [Figure S2].
@@
Figure 1: The room-temperature crystal structures and JT-ordering of (a) many JT-ordered perovskites such as KCuF3 and LaMnO3 Khomskii and Streltsov (2021) and (b) layered nickelates such as NaNiO2 and stoichiometric LiNiO2 Dyer et al. (1954); Phillips et al. (2025). Pink metal-oxygen bond lengths indicate JT-elongated bonds. Only octahedra associated with JT-active sites are displayed. In the perovskites, purple sites are the JT-distorted site (i.e. Mn, Cu, Cr) and green sites are the A-site (i.e. La, K). In the nickelates, grey sites are Ni3+ and yellow sites are the alkali metal.
cJT distortions also occur in the layered triangular-lattice nickelates NaNiO2 Dyer et al. (1954); Dick et al. (1997); Chappel et al. (2000); Sofin and Jansen (2005); Nagle-Cocco et al. (2022, 2024, 2025) and defect-free LiNiO2 Chung et al. (2005); Phillips et al. (2025). Most evidence indicates this is a collinear cJT distortion, where NiO6 octahedra within each layer all elongate along the same axis and this axis is preserved from layer to layer (shown in Figure 1(b)), although there is some evidence for zigzag-type intra-layer ordering in LiNiO2 Chung et al. (2005); Chen et al. (2011b); Foyevtsova et al. (2019) [Figure S3]. Conversely, some orbitally-degenerate nickelates such as AgNiO2 Wawrzyńska et al. (2007); Kang et al. (2007) or the nickelate perovskites García-Muñoz et al. (1994); Mizokawa et al. (2000); Garcia-Munoz et al. (2009) do not feature JT-distortions, but instead break degeneracy via charge disproportionation. There is some computational Chen et al. (2011a); Foyevtsova et al. (2019); Poletayev et al. (2025) and experimental Green et al. (2020); Takegami et al. (2024); Poletayev et al. (2025) evidence for a similar charge disproportionation mechanism, as an alternative to JT behaviour, occurring in LiNiO2.
@@
Beyond perovskites and layered nickelates, JT-driven orbital order occurs in a range of other materials. The K2PtCl6-type material (NO)2VCl6, which has a compressive, not elongational, distortion, exhibits a collinear cJT distortion within the $ab$-plane, but the axis of octahedral compression alternates with layer stacking Henke (2003). The spinels Mn3O4 Baron (1998), LiMn2O4 Yamaguchi et al. (1998), MgMn2O4 Yokozaki et al. (2021), and ZnMn2O4 Patra et al. (2019)[Figure S4], wolframite CuWO4 Schofield et al. (1997), Prussian Blue analogs such as CuPt(CN)6 Harbourne et al. (2024) and K2Cu[Fe(CN)6Cattermull et al. (2022), and Ruddlesden–Poppers like A2CuF4 (A=Na,K) Herdtweck and Babel (1981); Hiley et al. (2025), all exhibit simple collinear cJT distortions, although a pressure-induced change in cJT ordering has been observed in Na2CuF4 Hiley et al. (2025). The distorted trirutile CuSb2O6 Nakua et al. (1991) also exhibits intra-planar collinear ordering, but alternating $ab$-planes exhibit different axes of elongation. There are clearly a wide range of possible orbital orderings which may be energetically favourable, depending on the geometry of the octahedral interactions and the JT-active cation.
@@
Heating materials with a cJT distortion will often result in a symmetry-raising transition. In many perovskites, an order-disorder transition has been reported, in which JT distortions persist locally in octahedra even when cJT distortions are suppressed. For example, the canonical orbital-ordered material LaMnO3 shows the melting of long-range orbital order around $T_{\mathrm{JT}}\approx 750$ K Granado et al. (2000); Chatterji et al. (2003); Zhou and Goodenough (2003); Qiu et al. (2005); Kharlamova and Arulraj (2009); Chatterji et al. (2004); Souza et al. (2004); Mondal et al. (2011); Ramirez et al. (2011); Thygesen et al. (2017); Tran et al. (2019); Saint-Paul and Lejay (2004); Tragheim et al. (2025). Local structure probes such as neutron total scattering Qiu et al. (2005); Thygesen et al. (2017) and extended x-ray absorption fine structure (EXAFS) Souza et al. (2004) have typically been interpreted as showing that distortions persist locally within octahedra, and correlate over short lengthscales, although a recent study has indicated these experimental results may actually be consistent with a more complex high-temperature state without conventionally JT-distorted octahedra Batnaran et al. (2025). An order-disorder transition for Y1-xLaxVO3 has also been reported based on neutron total scattering Yano et al. (2014) for all values of $x$ in the range $0\leq x\leq 0.3$. KCrF3 shows a similar transition in crystal symmetry below $\sim$973 K Margadonna and Karotsis (2007), which is likewise proposed to be order-disorder Margadonna and Karotsis (2007) by analogy to LaMnO3. Surprisingly, the isostructural KCuF3 shows no sign of the melting of orbital order below its decomposition temperature Marshall et al. (2013).
@@
cJT distortions are generally suppressed due to the additional thermal energy in the system Gehring and Gehring (1975). Expansion of the crystal lattice weakens the effective interaction between local JT centers mediating cooperativity, reducing the enthalpy gain associated with orbital ordering. Increasing thermal energy also enables octahedra to dynamically cycle through axes of elongation corresponding to different local energy minima, which would suppress a time-averaged cJT distortion via a dynamic JT fluctuation Sicolo et al. (2020). Additionally, free energy becomes dominated by entropic contributions over enthalpic contributions on heating, and so enthalpy minimisation due to cJT will be less of a factor in determining the crystal structure Ahmed and Gehring (2005); this is important as there may be significant configurational entropy associated with orbital disorder depending on the geometry of the lattice in question Nagle-Cocco et al. (2024). An interpretation of JT transitions as order-disorder can therefore be explained in terms of maximising configurational entropy at high temperature.
@@
Figure 2: Diagram showing the JT configurations associated with zero and non-zero geometric energies as used in Equations 3 and 5. (a) The case for perovskites, which is a reproduction of the Potts model of Ahmed and Gehring Ahmed and Gehring (2005, 2006, 2009). (b) The case for intra-layer nearest-neighbour NiO6 octahedra within layered nickelates. Blue solid circles indicate JT-active sites, red solid circles (for the layered nickelate case) indicate O atoms, and solid lines indicate JT elongations.
@@
Layered nickelates LiNiO2 and NaNiO2 are promising battery materials Vassilaras et al. (2013); Steele et al. (2025); Bianchini et al. (2019); Genreith-Schriever et al. (2023), and so the nature of the JT transition in these materials has attracted much interest Nagle-Cocco et al. (2022, 2024); Radin et al. (2020); Nagle-Cocco et al. (2025); Genreith-Schriever et al. (2024); Phillips et al. (2025). Computational work based on Monte Carlo simulations integrated with density functional theory (DFT) Radin et al. (2020) concluded that although the parameter phase space for layered nickelates can accommodate both an order-disorder and a displacive transition, an order-disorder transition is more likely in common with previously-studied JT transitions. Radin et al. (2020) However, more recently, both experimental Nagle-Cocco et al. (2024); Jacquet et al. (2025) and ab initio molecular dynamics simulations Nagle-Cocco et al. (2024); Genreith-Schriever et al. (2024) have indicated that layered nickelates exceptionally exhibit displacive JT transitions, in which there is no minimum in the free energy landscape associated with JT distortions above the JT transition. This was attributed Nagle-Cocco et al. (2024) to the subextensive configurational entropy associated with orbital disorder on a 2D triangular lattice in which the axes of elongation are geometrically constrained by the chemical instability associated with having multiple axes of elongation pointing at a common O anion. In contrast, we note that evidence for an order-disorder transition in non-stoichiometric LiNiO2 with site-mixing has also been reported Jacquet et al. (2025).
@@
While the observation of displacive behaviour in nickelates appeared exceptional at the time they were reported, a recent study Batnaran et al. (2025) has shown that experimental data for the LaMnO3 case can be reproduced without well-defined and prevalent octahedral elongations. This would indicate that dynamics, rather than configurational entropy, is the key driver of behaviour through JT transitions. Further, it would indicate that displacive JT transitions may be more common, or even ubiquitous if a paradigm shift in our perspective JT transitions were to follow. This justifies close scrutiny of the evidence for previous assignments of order-disorder character to JT transitions.
@@
The cJT distortion in perovskites was previously studied in Monte Carlo simulations by Ahmed and Gehring (2005) Ahmed and Gehring (2005) and applied to the case of LaMnO3 Ahmed and Gehring (2006, 2009). In their model, an anisotropic Potts variable on each site successfully reproduced the cJT ordering at room-temperature in LaMnO3 Ahmed and Gehring (2005, 2006) and the dependence of lattice parameters on temperature through the JT transition Ahmed and Gehring (2009). However, their simple model did not include amplitudes of JT distortion or the possibility of JT-undistorted MnO6 octahedra, and so an order-disorder transition was implicitly built into the model. More recently, Tragheim et al. (2025) Tragheim et al. (2025) developed a more complex model in which JT amplitude may vary as a consequence of the octahedral tilts and shapes, which are the varying parameters; however, they do not apply this directly to the temperature-induced JT transition.
@@
For the layered nickelates, Radin et al. Radin et al. (2020) used an anharmonic vibrational Hamiltonian with continuous JT modes and Monte Carlo sampling, parametrized from DFT. Unlike the model of Ahmed and Gehring, this more complex approach Radin et al. (2020) contains a regime in the phase space of starting parameters, the stiff-lattice regime, which can lead to displacive behaviour. However, using parameters from DFT, they predicted that layered nickelates would occur in the soft-lattive regime where order-disorder transitions are predicted by the Monte Carlo simulations.
@@
In this present work, we extend the relatively simple model of Ahmed and Gehring with variable amplitude, $\rho$, of JT distortion on each site and a single-ion term in the Hamiltonian to simulate the enthalpy saving associated with JT distortion. Using this model, we simulate the JT transitions in perovskites such as LaMnO3. We then develop a very similar, simple model for the layered nickelates.
@@
We find that the mean magnitude of JT distortion decreases through the transition but remains non-zero, with the high-temperature magnitude being larger in the perovskite lattice than the 2D nickelate lattice, consistent with experimental results which are interpreted in terms of an order-disorder transition. However, the distribution of values of JT magnitude does not resemble that at low-temperatures. For an order-disorder transition, we would expect to see a well-defined peak in the distribution at finite, non-zero $\rho$, with negligible density at $\rho=0$. Instead, we see little preference for any particular $\rho$ value at high-temperatures, which appears consistent with dynamic exploration of the $E_{g}$($Q_{2}$$Q_{3}$) phase space. Our findings therefore support the reinterpretation of prior reports of order-disorder JT transitions first suggested by Batnaran et al. Batnaran et al. (2025), and suggest that displacive JT transitions may be more common than previously thought.

### v3

@@ Abstract
Jahn–Teller (JT) distortions are a key driver of physical properties in many correlated oxide materials. Cooperative JT distortions, in which long-range orbital order reduces the symmetry of the average structure, are common in JT-distorted materials at low temperatures. This long-range order will often melt on heating, via a transition to a high-temperature state without long-range orbital order. The nature of this transition has been observed to vary with different materials depending on crystal structure; in LaMnO3 the transition has generally been interpreted as order-disorder, whereas in layered nickelates AANiO2 (AA=Li,Na) there is a displacive transition. However, recent theoretical work has suggested that previous evidence for order-disorder may in fact be a consequence of phonon anharmonicity, rather than persistence of JT distortions. In this work, we run Monte Carlo simulations with a simple Hamiltonian that is modified to include terms dependent on the JT amplitude $\rho$, which is allowed to vary within the simulation via the Metropolis algorithm. Our simulations yield distributions of JT amplitudes consistent with displacive rather than order-disorder behaviour for both perovskites and layered nickelates, suggesting that displacive-like JT transitions may be more common than previously assumed. We also find significant differences between the transition observed for perovskites compared with layered nickelates, which we attribute to differing extensivity of configurational entropy on the two lattices, showing the crucial role of lattice geometry in determining behaviour.
@@ I Introduction
Transition metal oxide octahedra with electronic configurational degeneracy are susceptible to a Jahn–Teller (JT) distortion Jahn and Teller (1937); Öpik and Pryce (1957); Longuet-Higgins et al. (1958); Kanamori (1960); Gehring and Gehring (1975); Halcrow (2013); Goodenough (1998), causing their shape to deviate from that of a regular octahedron. Such distortions most commonly manifest as a linear superposition of the $E_{g}(Q_{2},Q_{3})$ van Vleck modes Van Vleck (1939); Nagle-Cocco and Dutton (2024), where $Q_{2}$ and $Q_{3}$ are planar rhombic and tetragonal distortions, respectively, although the tetragonal $Q_{3}$ component tends to dominate. JT distortions have been observed in both molecules Halcrow (2013) and crystals Goodenough (1998). The JT distortion is relevant to many phenomena including superconductivity in the cuprates Fil et al. (1992); Keller et al. (2008); Bussmann-Holder and Keller (2022), spin-orbit ordering Khomskii and Streltsov (2021), and ionic mobility Kim et al. (2015); Li et al. (2016), and can lead to structural transitions on cycling in battery electrode materials Kim et al. (2015); Li et al. (2016); Choi et al. (2019).
@@
Within solids, the axes of octahedral elongation can correlate via orbital ordering over macroscopic length-scales as a cooperative JT (cJT) distortion, typically manifesting as a decrease in unit cell symmetry Gehring and Gehring (1975); Goodenough (1998). For instance, the perovskites KMF3 (M=Cr,Cu) Zhou et al. (2011); Margadonna and Karotsis (2006), LaMnO3 Rodríguez-Carvajal et al. (1997), and [(CH3)2NH2]Cu(HCOO)3 Scatena et al. (2021) exhibit a cJT distortion with alternating MO6 (M=Cu,Cr,V,Mn) octahedra elongated along two orthogonal directions within the $xy$-plane, as shown in Figure 1(a). Such planes generally show C-type stacking along the $c$-direction, in which each plane has identical ordering Rodríguez-Carvajal et al. (1997); Ren et al. (1998); Blake et al. (2009); Zhou et al. (2011). Interestingly, similar orbital ordering driven by $t_{2g}$ orbitals is seen in the vanadate perovskites AVO3 (A=Ho,La,Y) Blake et al. (2009); Ren et al. (1998); Blake et al. (2001, 2002), although on cooling to low temperatures, some (such as LaVO3 and YVO3) have been shown to exhibit out-of-phase stacking along the $c$-direction Bordet et al. (1993); Blake et al. (2001, 2002) [SI Figure S1]. The flexibility of the perovskite structure means that other kinds of ordering can also be exhibited; for instance, Ba0.5La0.5CoO3 exhibits a collinear elongational cJT ordering of the CoO6 octahedra when the La/Ba is disordered Fauth et al. (2002) but a compressive collinear ordering when the La/Ba is ordered Nakajima et al. (2005) [Figure S2].
@@
Figure 1: The room-temperature crystal structures and JT-ordering of (a) many JT-ordered perovskites such as KCuF3 and LaMnO3 Khomskii and Streltsov (2021), corresponding to phase 5 of the phase diagram of Ahmed and Gehring Ahmed and Gehring (2005), and (b) layered nickelates such as NaNiO2 and stoichiometric LiNiO2 Dyer et al. (1954); Phillips et al. (2025). The stacking in (a) is commonly termed C-type orbital order. Pink metal-oxygen bond lengths indicate JT-elongated bonds. Only octahedra associated with JT-active sites are displayed. In the perovskites, purple sites are the JT-distorted site (i.e. Mn, Cu, Cr) and green sites are the A-site (i.e. La, K). In the nickelates, grey sites are Ni3+ and yellow sites are the alkali metal.
cJT distortions also occur in the layered triangular-lattice nickelate NaNiO2 Dyer et al. (1954); Dick et al. (1997); Chappel et al. (2000); Sofin and Jansen (2005); Nagle-Cocco et al. (2022, 2024, 2025), where diffraction shows a collinear ordering in which NiO6 octahedra within each layer elongate along a common axis that is preserved from layer to layer [Figure 1(b)]. By contrast, the electronic structure of LiNiO2 has long remained under debate, with the as-synthesised material typically not showing a comparable cJT distortion Dyer et al. (1954); Bianchi et al. (2001); Chung et al. (2005); Petit et al. (2006). This has been attributed to suppression of long-range orbital order by Ni/Li site mixing and related defects Petit et al. (2006); Lin et al. (2022); Genreith-Schriever et al. (2024); Phillips et al. (2025). Experimental evidence from neutron total scattering Chung et al. (2005) and X-ray absorption spectroscopy Rougier et al. (1995) shows evidence that JT distortion likely occurs locally in conventional LiNiO2, without long-range order. An alternative explanation is charge disproportionation, for which there is evidence from both computational studies Chen et al. (2011a); Foyevtsova et al. (2019); Poletayev et al. (2025) and X-ray absorption spectroscopy measurements Green et al. (2020); Takegami et al. (2024); Poletayev et al. (2025) by analogy to layered silver nickelate Wawrzyńska et al. (2007); Kang et al. (2007) and the nickelate perovskites García-Muñoz et al. (1994); Mizokawa et al. (2000); Garcia-Munoz et al. (2009). More recent studies of carefully prepared, stoichiometric LiNiO2 have reported evidence for a collinear cJT distortion similar to that of NaNiO2 Phillips et al. (2025).
@@
Beyond perovskites and layered nickelates, JT-driven orbital order occurs in a range of other materials. The K2PtCl6-type material (NO)2VCl6, which has a compressive, not elongational, distortion, exhibits a collinear cJT distortion within the $ab$-plane, but the axis of octahedral compression alternates with layer stacking Henke (2003). The spinels Mn3O4 Baron (1998), LiMn2O4 Yamaguchi et al. (1998), MgMn2O4 Yokozaki et al. (2021), and ZnMn2O4 Patra et al. (2019)[Figure S4], wolframite CuWO4 Schofield et al. (1997), Prussian Blue analogs such as CuPt(CN)6 Harbourne et al. (2024) and K2Cu[Fe(CN)6Cattermull et al. (2022), and Ruddlesden–Poppers like A2CuF4 (A=Na,K) Herdtweck and Babel (1981); Hiley et al. (2025), all exhibit simple collinear cJT distortions, although a pressure-induced change in cJT ordering has been observed the spinel in Na2CuF4 Hiley et al. (2025). The distorted trirutile CuSb2O6 Nakua et al. (1991) also exhibits intra-planar collinear ordering, but alternating $ab$-planes exhibit different axes of elongation. There are clearly a wide range of possible orbital orderings which may be energetically favourable, depending on the geometry of the octahedral interactions and the JT-active cation.
@@
Heating materials with a cJT distortion will often result in a symmetry-raising transition. In many perovskites, an order-disorder transition has been reported, in which JT distortions persist locally in octahedra even when cJT distortions are suppressed. For example, the canonical orbital-ordered material LaMnO3 shows the melting of long-range orbital order around $T_{\mathrm{JT}}\approx 750$ K Granado et al. (2000); Chatterji et al. (2003); Zhou and Goodenough (2003); Qiu et al. (2005); Kharlamova and Arulraj (2009); Chatterji et al. (2004); Souza et al. (2004); Mondal et al. (2011); Ramirez et al. (2011); Thygesen et al. (2017); Tran et al. (2019); Saint-Paul and Lejay (2004); Tragheim et al. (2025). Local structure probes such as neutron total scattering Qiu et al. (2005); Thygesen et al. (2017) and extended X-ray absorption fine structure (EXAFS) Souza et al. (2004) have typically been interpreted as showing that distortions persist locally within octahedra, and correlate over short lengthscales. KCrF3 shows a similar transition in crystal symmetry below $\sim$973 K Margadonna and Karotsis (2007), which is likewise proposed to be order-disorder Margadonna and Karotsis (2007) by analogy to LaMnO3. Surprisingly, the isostructural KCuF3 shows no sign of the melting of orbital order below its decomposition temperature Marshall et al. (2013).
@@
cJT distortions are generally suppressed due to the additional thermal energy in the system Gehring and Gehring (1975). Expansion of the crystal lattice weakens the effective interaction between local JT centers mediating cooperativity, reducing the enthalpy gain associated with orbital ordering. Increasing thermal energy also enables octahedra to dynamically cycle through axes of elongation corresponding to different local energy minima, which would suppress a time-averaged cJT distortion via a dynamic JT fluctuation Sicolo et al. (2020). Additionally, free energy becomes dominated by entropic contributions over enthalpic contributions on heating, and so enthalpy minimisation due to cJT will be less of a factor in determining the crystal structure Ahmed and Gehring (2005); this is important as there may be significant configurational entropy associated with orbital disorder depending on the geometry of the lattice in question Nagle-Cocco et al. (2024). The common interpretation of JT transitions as order-disorder can therefore be explained in terms of maximising configurational entropy at high temperature.
@@
Figure 2: Diagram showing the JT configurations associated with zero and non-zero geometric energies as used in Equations 3 and 5. (a) The case for perovskites, which is a reproduction of the Potts model of Ahmed and Gehring Ahmed and Gehring (2005, 2006, 2009). (b) The case for intra-layer nearest-neighbour NiO6 octahedra within layered nickelates assuming all sites have $\rho=1$. Blue solid circles indicate JT-active sites, red solid circles (for the layered nickelate case) indicate O atoms, and solid lines indicate JT elongations.
@@
The nature of the JT transition in the previously-discussed layered nickelate materials has attracted much interest Nagle-Cocco et al. (2022, 2024); Radin et al. (2020); Nagle-Cocco et al. (2025); Genreith-Schriever et al. (2024); Phillips et al. (2025). Computational work based on Monte Carlo simulations integrated with density functional theory (DFT) Radin et al. (2020) concluded that although the parameter phase space for layered nickelates can accommodate both an order-disorder and a displacive transition, an order-disorder transition is more likely in common with previously-studied JT transitions. Radin et al. (2020) However, more recently, X-ray absorption spectroscopy Nagle-Cocco et al. (2024); Jacquet et al. (2025), neutron total scattering Nagle-Cocco et al. (2024), and ab initio molecular dynamics simulations Nagle-Cocco et al. (2024); Genreith-Schriever et al. (2024) have indicated that layered nickelates exceptionally exhibit displacive JT transitions, in which there is no minimum in the free energy landscape associated with JT distortions above the transition temperature. This was attributed Nagle-Cocco et al. (2024) to the subextensive configurational entropy associated with orbital disorder on a 2D triangular lattice in which the axes of elongation are geometrically constrained by the chemical instability associated with having multiple axes of elongation pointing at a common O anion. In contrast, we note that evidence for an order-disorder transition in non-stoichiometric LiNiO2 with site-mixing has also been reported from X-ray absorption spectroscopy measurements Jacquet et al. (2025).
@@
While the observation of displacive behaviour in nickelates appeared exceptional at the time they were reported, a recent study Batnaran et al. (2025) has shown that experimental total scattering data for the LaMnO3 case can be reproduced without well-defined and prevalent octahedral elongations. This would indicate that dynamics, rather than configurational entropy, is the key driver of behaviour through JT transitions. This raises the possibility that displacive JT transitions are more common than previously assumed, and that many historical assignments of order-disorder character may warrant re-examination.
@@
The cJT distortion in perovskites was previously studied using Monte Carlo simulations by Ahmed and Gehring (2005) Ahmed and Gehring (2005) and applied to the case of LaMnO3 Ahmed and Gehring (2006, 2009). In their model, an anisotropic Potts variable on each site successfully reproduced the cJT ordering at room-temperature in LaMnO3 Ahmed and Gehring (2005, 2006) and the dependence of lattice parameters on temperature through the JT transition Ahmed and Gehring (2009). However, their simple model did not include amplitudes of JT distortion or the possibility of JT-undistorted MnO6 octahedra, and so an order-disorder transition was implicitly built into the model. More recently, Tragheim et al. (2025) Tragheim et al. (2025) developed a more complex model in which JT amplitude may vary as a consequence of the octahedral tilts and shapes, which are the varying parameters; however, they did not apply this directly to the temperature-induced JT transition.
@@
For the layered nickelates, Radin et al. Radin et al. (2020) used an anharmonic vibrational Hamiltonian with continuous JT modes and Monte Carlo sampling, parametrized from DFT. Unlike the model of Ahmed and Gehring, this more complex approach Radin et al. (2020) contains a regime in the phase space of starting parameters, termed the “stiff-lattice regime”, which can lead to displacive behaviour. However, they predicted that layered nickelates would occur in the soft-lattice regime where order-disorder transitions are predicted by the Monte Carlo simulations.
@@
In this present work, we extend the relatively simple model of Ahmed and Gehring with variable amplitude, $\rho$, of JT distortion on each site and a single-ion term in the Hamiltonian to simulate the enthalpy saving associated with JT distortion. Using this model, we simulate the JT transitions in perovskites such as LaMnO3. We then develop a similar, simple model for the layered nickelates.
@@
We find that the mean magnitude of JT distortion decreases through the transition but remains non-zero, with the high-temperature magnitude being larger in the perovskite lattice than the 2D nickelate lattice, consistent with experimental results which are interpreted in terms of an order-disorder transition. However, the distribution of values of JT magnitude does not resemble that at low temperatures. For an order-disorder transition, we would expect to see a well-defined peak in the distribution at finite, non-zero $\rho$, with negligible density at $\rho=0$. Instead, we see little preference for any particular $\rho$ value at high temperatures for perovskites, which appears consistent with dynamic exploration of the $E_{g}$($Q_{2}$$Q_{3}$) phase space. Our findings therefore support the reinterpretation of prior reports of order-disorder JT transitions first suggested by Batnaran et al. Batnaran et al. (2025), and suggest that displacive JT transitions may be more common than previously thought.